OpenStudy (anonymous):

the sum of the digits of a three digit is 11. If the first and last digit are interchanged, the new number is greater than the original number by 594. If the tens and the units digit are interchanged, the resulting number is greater than the original number by 36. Find the Original number.

OpenStudy (anonymous):

Please help me..

OpenStudy (anonymous):

\[a+b+c=11\\ 100c+10b+a=100a+10b+c=594\\ 100a+10c+b=100a+10b+c+36\]